### v1

# What Did My Car Say? Autonomous Vehicle Explanation Errors, Context, and Personal Traits Impact Comfort, Reliance, Satisfaction, and Driving Confidence

Abstract.
Explanations for autonomous vehicle (AV) decisions may build trust, however, explanations can contain errors. In a simulated driving study (n = 232), we tested how AV explanation errors, driving context characteristics (perceived harm and driving difficulty), and personal traits (prior trust and expertise) affected a passenger’s comfort in relying on an AV, preference for control, confidence in the AV’s ability, and explanation satisfaction. Errors negatively affected all outcomes. Surprisingly, despite identical driving, explanation errors reduced ratings of the AV’s driving ability. Severity and potential harm amplified the negative impact of errors. Contextual harm and driving difficulty directly impacted outcome ratings and influenced the relationship between errors and outcomes. Prior trust and expertise were positively associated with outcome ratings. Results emphasize the need for accurate, contextually adaptive, and personalized AV explanations to foster trust, reliance, satisfaction, and confidence. We conclude with design, research, and deployment recommendations for trustworthy AV explanation systems.
Autonomous Vehicles, Explainable AI, AI Errors
ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studies
1. Introduction
Autonomous vehicles offer a wide range of potential societal benefits, including reduced driving infractions, traffic volume, environmental impact, and passenger stress (Fagnant and Kockelman, 2015). Despite these benefits, it is a well-documented problem that adoption is limited by a lack of trust in how vehicles make decisions (Kenesei et al., 2022). This is not a problem specifically with autonomous vehicles, but a widespread issue across many types of AI-based systems (Bedué and Fritzsche, 2022). System transparency via explainable AI (XAI) has been proposed as a means to mitigate concerns with AI-based systems like AVs, offering users a look “under the hood” of black-box AI models so they understand what the system is doing and why  (Gunning and Aha, 2019; Miller, 2019).
However, although potentially increasing trust, explanation of AI interactions in the real world can contain errors. Prior work has shown that when autonomous vehicles exhibit driving errors, passenger trust and willingness to rely on the vehicle can deteriorate quickly (Seet et al., 2020; Kaplan et al., 2023). Far less is known about the impact of explanation errors (Cabitza et al., 2024; Kenny et al., 2021). Particularly for safety-critical systems like autonomous vehicles – that may rely on explanations and other in-vehicle communications to elicit trust comfort, and safe reliance with users – knowing the consequences of errors is pivotal to safe deployment. Understanding these consequences is essential for user reliance because, without this knowledge, AV systems cannot be deployed safely or ethically (Martinho et al., 2021).
In addition, human-AV interactions do not exist in a vacuum: they are sensitive to the contextual demands of the external driving environment (Hoff and Bashir, 2015). It is important to consider the driving context in which explanation errors occur, as prior work shows that context may significantly influence how people interact with AI-based systems (Lim et al., 2009; Schilit et al., 1994), including autonomous vehicles (Capallera et al., 2022; De Salis et al., 2020). The complexity of a driving situation and perceived risk of harm have been singled out as particular factors of interest (Ha et al., 2020; Kaufman et al., 2024b), as these might drive a person’s explanatory needs and reliance behaviors.
Finally, recent work has posited that AV explanations should be tailored to meet the specific needs of the people interacting with the system (Ma and Feng, 2023; Kaufman et al., 2024b). It is well known that people may interact differently with AI-based systems (Schneider and Handali, 2019) based on the individual characteristics or prior experiences they may have (Ayoub et al., 2021b). Particularly well studied are domain expertise (Araujo et al., 2020; Pazzani et al., 2022; Kaufman and Kirsh, 2022) and initial (dispositional) trust (Hoff and Bashir, 2015).
In the present study, we examine the impact of explanation errors across a variety of realistic, simulated driving scenarios. To deepen our investigation and understand the significance of the type of error presented, participants were shown AV explanations at three distinct accuracy levels. We measured the effect of explanation errors on four main outcomes: comfort relying on the AV, preference for control, confidence in the AV’s driving ability, and explanation satisfaction. We include measures of scenario context – perceived harm and driving difficulty – to assess their impact on our driving outcomes, including how they may moderate the relationship between explanation errors and user perceptions. To explore the influence of individual differences, we also included measures of trust and expertise to determine if they predict study outcomes.
Our findings reveal that explanation errors, contextual characteristics, and personal traits significantly impact how a person may think, feel, and behave towards AVs. Explanation errors negatively affected all outcomes, with impacts proportional to the magnitude and negative implications of the error. Harm and driving difficulty directly impacted outcomes as well as moderated the relationship between errors and outcomes, though in opposing ways. Overall, harm was generally seen as more important and more negative than difficulty. Participants with higher expertise tended to trust AVs more, and these each correlated with more positive outcome ratings in turn.
In sum, we contribute:
  • A study on the impact of explanation errors based on information type (‘why’ and ’what + why’).
  • An assessment of how contextual harm and driving difficulty, initial trust, and expertise impact study outcomes.
  • A set of future design and research implications based on the present findings.
Results highlight the critical need for accurate and contextually adaptive explanations for autonomous vehicles to enhance user trust, reliance, satisfaction, and confidence. Recognizing the implications of explanation errors is vital for advancing AV research and guiding design teams to make informed decisions. This understanding lays the groundwork for developing context-aware designs, personalized explanation interfaces, and establishing ethical or regulatory guidelines to ensure the deployment of safe and trustworthy explainable AI (XAI) systems for autonomous vehicles.
2. Related Work
2.1. AV Trust and Explainability
Human-Centered Explainable AI – Continuous development of explainable AI (XAI) has brought major progress in the quest for trust through AI transparency (Miller, 2019). Approaches to XAI vary, often limited by the availability of the model (Simonyan et al., 2013; Ribeiro et al., 2016). Recent work found that – even when an explanation is given – trust and engagement with the system may not improve unless the right information is given in the right way, at the right time (Wang et al., 2019; Liao and Varshney, 2021). Several studies and theory pieces have demonstrated the value of human-interpretable explanations on user understanding and trust  (Holzinger et al., 2019; Soltani et al., 2022). In particular, explanations that are modeled off those given by human experts have suggested as a means to increase understanding for experts and novices alike (Pazzani et al., 2022; Kaufman and Kirsh, 2022). There have likewise been calls for the need for explanations to be sensitive to particular user characteristics (such as personal traits or experiences)  (Ehsan and Riedl, 2020; Ehsan et al., 2021; Kaufman and Kirsh, 2023) or context of use (including the specific use environment and goals of a user) (Schilit et al., 1994; Kaufman et al., 2024b). Recent discussion on XAI usefulness had proposed that explanations should move beyond simply conveying algorithmic mechanisms for decision-making and instead supporting user hypothesis generation (Miller, 2023). Despite the large amount of research on explanation design, little has been done to understand what happens when these explanations fail. The study we present here is a first step towards filling this knowledge gap, using autonomous vehicles as a specific domain of interest.
Interface Modalities – Research on in-vehicle interfaces for explanation, such as visual heads-up-displays (HUDs) (Currano et al., 2021; Chang et al., 2016; Schartmüller et al., 2019), audio interaction (Mok et al., 2015; Jeon et al., 2009; Löcken et al., 2017), and even haptic feedback for drivers (Di Campli San Vito et al., 2020) has shown that there are a variety of ways explanations can support users, each dependent on the use-case and context of interest. The benefits and drawbacks of different modalities often relate to the complex process of transferring sufficient knowledge to accomplish a task (such as to build understanding through system transparency) without creating too much cognitive load (Kaufman et al., 2024a; Kim et al., 2023; Colley et al., 2021). In the present study, we leverage video and audio explanations for our study on the impact of errors, as these are common and efficient methods to transfer information to a user without overwhelm. We present both modalities of information at the same time to increase the accessibility of our study and ensure there is successful transfer of information.
Explanation Content – Choosing the appropriate content for an AI explanation is crucial for enhancing rider trust and reliance (Miller, 2019). Recent work as focused on providing a description of what a vehicle is doing and why a vehicle is doing it, as these enable a user to create a momentary evaluation of the AV’s behavior in terms of reliance (Hoff and Bashir, 2015). For example, Koo et al. (2015) present ’how’, ‘why’, and a combination of both in various simulated autonomous driving scenarios to assess driver attitudes and safety performance. They found improved safety when both ’how’ and ‘why’ were presented to drivers, but a preference for ‘why’ explanations alone. Kaufman et al. (2024a) leveraged auditory and visual explanations to teach humans to be better drivers via an ’AI coach’, highlighting the additive value of what and why-type information for transferring knowledge, adding that too much information can cause overwhelm. They emphasize the need for explanations to strike a balance between complexity and comprehensiveness. The impact of what (similar to Koo’s how) and why-type explanation errors – explored in the present study – remains unknown.
Factors Impacting Trust – Knowing what factors may impact a person’s trust and reliance decisions is vitally important to designing XAI explanations to support them. Hoff and Bashir’s theoretical model of trust in automation highlights the importance of three distinct yet interdependent facets of trust: dispositional (e.g. personality or cultural attitudes), situational (e.g. based on a particular context of use), and learned trust (e.g. based on a present evaluation of system performance) (Hoff and Bashir, 2015). Additionally, Kaufman et al. (2024b) developed a framework to understand situational awareness in joint action between humans and AV, a key focus of explainable AI transparency in safety-critical situations like driving. They describe how communications like AV explanations can enable human-AV teamwork to achieve particular goals like safe and trustworthy driving. Factors of interest include external driving conditions, human traits and abilities, and communication preferences and goals – all of which are crucial in managing driving difficulty and reducing the risk of harm. Using these system-based models as a backdrop, AI explanation designers can form hypotheses on how to build more trustworthy systems. In the present study, we build off these models by investigating specific driving context factors and personal traits which may impact a person’s reliance judgments and, in the case of context, how much an error matters.
Knowledge Gap – Indeed, explanation interfaces have been implemented in mainstream deployed autonomous vehicles, such as communication interfaces by Tesla (Inc., [n. d.]) and Waymo (LLC, [n. d.]). Despite these efforts, we still know very little on the impact of AV explanation errors on how a person will trust and interact with an AV. We know even less about how errors may depend on contextual factors like driving difficulty or harm (Capallera et al., 2022; De Salis et al., 2020; Ha et al., 2020). With this work, we seek to understand how contextual harm and difficulty of driving may affect study outcomes, as well as contribute to the body of literature on how personal traits predict a person’s interactions with an AV.
2.2. AI Errors
Widespread integration of AI systems into everyday tasks has demonstrated huge benefits to productivity and optimization in many domains (Fauzi et al., 2023). However, concerns over AI errors remain a major point of contention, particularly as systems proliferate. Examples of errors with real-world consequences include language models’ propensity to hallucinate (Xu et al., 2024) and algorithmic bias that favors men over women used in the hiring process (Dastin, 2018).
General AV Errors – Autonomous vehicle driving errors have important real-world consequences, which may include physical harm or even fatalities (Favarò et al., 2017). Even in cases where AV driving outperforms humans (Schwall et al., 2020), concerns over vehicle errors can be a major hindrance to adoption and use (Choi and Ji, 2015). Luo et al. (2020) shows that errors caused by an AV had a more significant negative impact on user trust than external errors, such as those caused by other drivers or road conditions. Declines in trust may be difficult to recover from and have long-lasting effect (Seet et al., 2020). Explanations are no panacea: trust is difficult to achieve when the system itself performs poorly, even when explanations meant to elicit trust are presented (Kaplan et al., 2023). In this paper, we seek to connect prior work showing the dire impact of driving errors (Luo et al., 2020; Zhang et al., 2022) to broader XAI literature investigating user perceptions of AI explanations (Lebovitz, 2019; Cabitza et al., 2024).
Explanation Errors – Some prior work has investigated the impact of explanation errors for autonomous systems. In a study on “white box” XAI, Cabitza et al. (2024) found that non-expert users tend to not catch explanations errors and believe the system even when explanations were wrong, attributing the phenomena to the Halo Effect found in social psychology where people assumed correctness of the system without verifying accuracy. Over- or under-reliance is a problem as people learn to calibrate their interactions with AI systems (Endsley, 2018). Conflicting results suggest that explanations may help reduce over-reliance in some cases (Vasconcelos et al., 2023), but increase over-reliance in others (Kenny et al., 2021). Other research has shown that the influence of explanations on reliance may be based on the systems’ performance itself (Papenmeier et al., 2022).
Knowledge Gap – Though several prior studies have investigated the consequences of accurate AV explanations, the impact of explanation errors on reliance behavior and related outcomes remains unexplored. Of particular interest to us are cases when the autonomous vehicle’s driving performs properly, as this allows us to separate the impact of driving performance from the impact of explanation performance. This has yest to be explored, and we address this important knowledge gap.
3. Method
We conducted an online experiment using realistic, simulated driving scenarios of an autonomous vehicle driving through various environments. These ranged from rural to urban driving and from routine navigational challenges like driving around construction cones to challenging situations like avoiding collisions with erratic drivers. During the experiment, participants viewed videos of the scenarios and, after each video, provided ratings related to trust, reliance, satisfaction, and evaluation of the AV.
To test the impact of explanation errors, participants were shown three versions of each scenario. Those with an “accurate” explanation were correctly told ‘what’ the AV was doing and ‘why’ it was doing it. Errors were introduced via a “low” error condition, where the AV correctly explained ‘what’ it was doing but incorrectly explained ‘why’, and a “high” error condition, where both ‘what’ and ‘why’ explanations were incorrect. In all three versions of each scenario, the AV drove identically, so only the explanations differed. The AV drove accurately and lawfully in all cases. Participants were randomly shown videos from 9 of 27 possible scenarios, providing ratings after each video, making 27 total video ratings per participant (9 scenarios X 3 accuracy levels). Scenarios were presented in random order. The 27 possible scenarios are described in Table 1.
Table 1. Scenario Descriptions
Scenario Description
10s1 Left turn in busy intersection, navigating around object in road.
10s2 Slowing to avoid collision between two vehicles ahead of the ego vehicle.
10s4 Parallel parking between two cars on side of busy road.
10s5 Slowing to avoid being sideswiped during turn by a vehicle that crosses into ego vehicle’s lane.
10s6 Sudden stop mid-intersection by large lead vehicle in adverse weather.
10s7 Left turn quickly followed by right merge for quick right turn in urban environment in adverse weather.
10s8 Ego vehicle slows for pedestrian crossing road in non-crosswalk area.
4s1 Reversing in parking lot.
4s2 Overtaking cyclist in suburban environment.
4s3 Ego vehicle must stop quickly for a pedestrian who jumps into the road.
4s4 Merging from far right lane to far left lane on highway to avoid emergency vehicles / car accident.
4s5 Ego vehicle must avoid object that falls off of lead vehicle on highway at night.
4s6 Ego vehicle hydroplanes on wet road on highway, and needs to maintain control.
4s7 Ego vehicle needs to pull over for flat tire at high speed.
4s8 Ego vehicle merges left to enter a busy highway.
5s2 Ego vehicle makes blind turn in intersection due to obstructed view. [removed]
5s3 Ego vehicle turning left and avoids a collision with another vehicle who ran a red light (T-bone).
5s4 Car in parallel lane merges into ego vehicle’s path.
5s5 Ego vehicle needs to navigate around a stopped cyclist. [removed]
5s6 Ego vehicle navigates construction zone at night.
7s1 Hidden stop sign at night.
7s3 Lead vehicle quickly decelerates (brake check)
7s4 Ego vehicle crosses the midline to overtake a slow lead vehicle.
7s5 Ego vehicle needs to slow quickly from high speed for a stopped cyclist around a turn. [removed]
7s6 Vehicle failure (flat tire) in small parking lot.
xs2 Ego vehicle waits for child to cross crosswalk before turning right at stop sign.
xs3 Ego vehicle stops during right turn to avoid collision with vehicle turning right from incorrect (left) lane.
We evaluated changes across four major outcomes of interest, with two additional descriptive outcomes, making a total of six ratings per scenario video. After each video, participants rated their: (1) comfort relying on the AV, (2) preference to take control, (3) satisfaction with the explanation, and (4) confidence in the AV’s driving ability. We hypothesized that ratings may be context-dependent. As such, we also collected two ratings describing the driving context: (5) perceived harm and (6) perceived difficulty of driving in each scenario. These context descriptor variables were collected after the accurate explanation videos only. Outside of the rating task, participants answered questions about their trust in AVs, expertise, and demographics.
Figure 1. Example images from four driving scenarios: (1) the AV slows for a pedestrian crossing the road on a foggy day in the city, (2) the AV is cut off by a vehicle turning right from the left lane, (3) the AV merges around a slow lead vehicle in a forested area, (4) the AV moves past a cyclist on a suburban street. In the bottom right corner of each scenario, written AV explanations appear on the vehicle’s dashboard display.
[…]
Example images from four driving scenarios: (1) the AV slows for a pedestrian crossing the road on a foggy day in the city, (2) the AV is cut off by a vehicle turning right from the left lane, (3) the AV merges around a slow lead vehicle in a forested area, (4) the AV moves past a cyclist on a suburban street. In the bottom right corner of each scenario, written AV explanations appear on the vehicle’s dashboard display.
3.1. Participants
A total of 232 participants spread throughout the United States participated in the study and completed all study procedures. Participants were recruited from the general population via existing participant email lists and via SONA, [^1] https://www.sona-systems.com an undergraduate study pool system where students are granted study credit for participation. The mean age of the study sample was 23.6 (SD = 11.3), with ages ranging from 18 to 85 years. The sample was 73.7% female.
3.2. Simulated Driving Scenarios
All simulated driving videos were custom-made by the research team using DReye VR (Silvera et al., 2022), a tool for creating realistic driving scenarios using the open-source driving simulator CARLA (Dosovitskiy et al., 2017). Scenario design was inspired by the National Highway Traffic Safety Administration (NHTSA) list of common driving situations that result in vehicle crashes (Najm et al., 2007). Examples include unexpected pedestrians in the road, collision avoidance from erratic other drivers, construction and emergency vehicle zones, parallel parking and reversing from park, navigating around stopped or slowed vehicles or cyclists, and dealing with flat tires or hydroplaning. Videos lasted between 10 and 30 seconds each (Examples in Fig. 1).
Data from a total of 27 scenarios (see Table 1) was collected during the study; three scenarios were removed during analysis due to the AV’s driving determined to be imperfect (going above the speed limit, crossing over the center line during a turn, and improper yielding), making a total of 24 included in the final analysis. Data was cleaned prior to analysis to ensure that all videos were viewed in full.
3.3. AV Explanations and Errors
During all driving scenario videos, explanations of the AV’s behavior were provided to participants at the time of AV action. Explanations were presented simultaneously in visual and auditory form to reflect current trends in AV explanation research (Kaufman et al., 2024a), as well as to provide better accessibility to participants. Specifically, explanations were presented visually in written English on the vehicle’s dashboard display and in auditory fashion via spoken English produced by Amazon Polly Text-To-Speech (Inc., 2023). For an example of the visual presentation of explanations, see Figure 1.
Explanations provide information on ‘what’ action the AV is doing (e.g. braking) as well as a local explanation for ‘why’ the vehicle is doing it (e.g. to avoid a collision with a merging vehicle). The importance of ‘what’ and ‘why’-type information has been studied in past experimental work on explanations of driving behavior by Kaufman et al. (2024a) and Koo et al. (2015). Frameworks by Wang et al. (2019) and Lim et al. (2019) in cognitive science and Miller (2019) in philosophy emphasize that both ‘what’ and ‘why’ information are important for transparency, trust, and user understanding when interacting with AI-based systems like AVs. The work presented here examines the impact of both ‘why’ errors alone, and ’why and what’ errors combined, respectively.
Explanation errors were presented via three conditions differing in the accuracy of the explanation provided. Table 2 shows a breakdown of the conditions and provides an example taken from one of the scenarios used in the study. Comparing the “accurate” to “low” group shows the impact of errors related to the AV’s rationale for behavior (i.e. ‘why’). Comparing the “low” to “high” group shows the impact of adding errors related to the AV’s description of its own action (i.e. adding ‘what’ to the ‘why’ error). It is important to note that the AV’s actual driving behavior remained consistent, accurate, and lawful across conditions; only the explanation itself changed. Participants rated videos for 9 of 27 possible scenarios – seeing one video for each of the three conditions per scenario – presented in random order. This makes a total of 27 total video ratings per participant (9 scenarios X 3 accuracy levels).
As part of a secondary analysis, we further categorized mistakes in the “high” error condition – where the AV is providing an incorrect description of what it is doing – based on the potential harm that would result should the AV have acted on the mistaken ‘what’ description. For example, if the AV mistakenly says it is going “left” when really it is going straight, we hypothesized that the impact of this error would be greater if going left would result in an accident, as opposed to simply a wrong turn. To explore if this difference in pragmatics does impact our results, we run a secondary analysis within just the ’high’ condition to test for impact of potential harm.
Table 2. Experimental Conditions. Participants saw videos across each condition, providing ratings for each.
Error Condition Explanation Description Example
Accurate Correct ‘what’ and correct ‘why’ “Braking, a pedestrian is crossing the road.”
Low Correct ‘what’ and incorrect ‘why’ “Braking, a cyclist is crossing the road.”      (When the obstacle is actually a pedestrian)
High Incorrect ‘what’ and incorrect ‘why’ “Merging right, a cyclist is crossing the road.” (When the vehicle is slowing, not merging, and the obstacle is actually a pedestrian)
3.4. Measures
3.4.1. Main Outcomes: Scenario Ratings
These measures were taken after each video to provide insight into the impact of the explanation errors.
  • Comfort relying on AV (proxy for trust). “How comfortable would you feel relying on this AV in this specific situation?” (0-10 rating scale)
  • Reliance on AV (preference to take control). “If this specific situation were to happen in the real world, would you prefer to rely on an AV or take control yourself?” (Binary choice: ’Rely on AV’ or ’Take control myself’). To aid comparison, reliance data was transformed from 0-1 to 0-10 to be on the same scale as the other variables. This does not impact how the data is interpreted.
  • Satisfaction with explanation. “How satisfied are you with the AV’s explanation?” (0-10 rating scale)
  • Confidence in AV driving ability. “Please rate your confidence in the AV’s driving ability.” (0-10 rating scale). Note: the actual driving performance never changed.
3.4.2. Context Descriptors
These measures were taken for each scenario after the ”accurate” condition video only to provide insight into the impact of driving context on main outcomes.
  • Harm of Driving Situation. “In the real world, how would you rate the risk of harm of this specific driving situation?” (0-10 rating scale)
  • Difficulty of Driving. “In the real world, how would you rate the difficulty of driving in this specific situation?” (0-10 rating scale)
3.4.3. Additional Outcomes
Additional outcomes further contextualize our main experimental findings. These were measured before and/or after the rating task.
  • Expertise. Expertise was measured before the rating task via three self-rated questions related to a person’s knowledge and understanding of AVs. (5-point likert scale from Strongly Disagree to Strongly Agree)
  • Trust. Trust was measured before and after the rating task via four questions related to adaptability, safety, overt trust, and willing to recommend a friend to ride in an AV. (5-point likert scale from Strongly Disagree to Strongly Agree)
  • Explicit Factors Contributing to Reliance Decisions. After the task, participants rated the relative impact of seven aspects of the driving and explanation experience on their reliance decision-making. (5-point likert scale from Not At All to Very Much). Factors included: potential harm, driving difficulty, accuracy of ‘what’ and ‘why’ information, AV driving performance, prior knowledge, and previous scenarios viewed.
4. Results
4.1. Impact of Explanation Errors: Comfort, Reliance, Satisfaction, and Confidence
4.1.1. Summary of Main Outcomes
Across our four main outcomes, segmenting the data by explanation error condition level gives us an initial impression on the impact of our experimental manipulation. We find the highest scores for comfort relying on the AV, reliance preference, satisfaction with the AV’s explanation, and confidence in the AV’s driving ability in the Accurate condition, followed by the Low condition and then the High condition (Table 3). Visualizing main outcomes by scenario, we find that the effect of condition was very consistent across scenarios (Figure 2.
Even when explanations were accurate, participants’ overall comfort relying on the AV (comfort), and their preference to rely on the AV (reliance), were middling to low, reflecting the overall reluctance to trust AVs found in past human-AV interaction research by Kenesei et al. (2022). Participants were more positive about the AV explanations provided to them in the study, however, this satisfaction deteriorated quickly when errors were introduced. Despite the AV’s driving performance – and therefore, demonstrated ability – remaining consistent across all conditions, impressions of the AV’s driving ability worsened as AV explanation errors were introduced.
Figure 2. Mean Outcome by Error Condition Across All Scenarios. Plot A shows mean scores for Comfort Relying on AV, Plot B shows mean scores for Reliance Preference (Binary, scaled to 1-10), Plot C shows mean scores for Satisfaction with AV Explanation, and Plot D shows mean scores for Confidence in AV Driving Ability. Results consistently demonstrated the trend of Accurate ¿ Low ¿ High.
[…]
Four visualizations, one for each major outcomes by error condition, shown for all 24 scenarios, depicting the mean outcome score by error condition across. Plot A shows mean scores for Comfort Relying on AV, Plot B shows mean scores for Reliance Preference (Binary, scaled to 1-10), Plot C shows mean scores for Satisfaction with AV Explanation, and Plot D shows mean scores for Confidence in AV Driving Ability. Results consistently demonstrated the trend of Accurate ¿ Low ¿ High.
Table 3. Means and Standard Errors of Main Outcomes By Error Condition. As errors increased, outcome ratings decreased.
Accurate Low High
Comfort Relying on AV 4.59 (0.06) 3.48 (0.06) 2.71 (0.06)
Reliance Decision 3.65 (0.11) 2.43 (0.10) 1.60 (0.09)
Satisfaction w/ Expl. 6.38 (0.06) 3.13 (0.06) 2.20 (0.06)
Confidence in Driving 5.16 (0.06) 3.68 (0.06) 2.84 (0.06)
4.1.2. Overall Impact of Errors on Main Outcomes
Linear mixed-effects (LME) models were used to measure the effect of errors on each major outcome: comfort relying on the AV, reliance decision, satisfaction with explanation, and confidence in the AV’s driving ability. Though LME models produce similar results to mixed-model ANOVAs, they offer greater flexibility for repeated measures experiments. They also incorporate random effects, which help reduce the likelihood of a Type 1 error (Boisgontier and Cheval, 2016). Models were made with fixed effects for error condition and random effects for the specific scenario and individual participant. The fixed effects allow us to compare differences between study conditions, while the random effects allow us to control for differences by scenario and individual participant.
Individual comparisons between each condition (Accurate-Low, Accurate-High, Low-High) show highly significant effects (p ¡ 0.001) across all four outcome measures (Figure 4), implying error condition significantly impacted outcomes. The results for comfort relying on the AV, reliance decision, and satisfaction with the explanation follow the expected trend: we find that outcome scores decrease as error level increases. Surprisingly, we found the same effect on a person’s evaluation of the AV’s driving ability, where confidence scores also decrease as error level increases. This is unexpected, given that the driving performance shown in the videos were identical and only the explanations changed. The implication is that there is a cross-over effect between a person’s evaluation of the explanation and the person’s evaluation of the vehicle’s driving, despite evidence suggesting that the driving performance is consistently high quality.
Table 4. Comparing Error Conditions for Main Outcomes (LME Models). These show highly significant outcome score differences between each error level.
Acc.-Low Acc.-High Low-High
$\beta$ df t sig. $\beta$ df t sig. $\beta$ df t sig.
Comfort Relying on AV -1.08 5114 -17.0 *** -1.85 5115 -28.9 *** -0.75 5110 -11.9 ***
Reliance Decision -1.19 5114 -10.2 *** -2.00 5115 -17.1 *** -0.80 5108 -6.8 ***
Satisfaction w/ Expl. -3.24 5116 -45.0 *** -4.16 5118 -57.8 *** -0.93 5111 -12.8 ***
Confidence in Driving -1.46 5114 -25.5 *** -2.29 5114 -38.3 *** -0.83 5111 -13.8 ***
Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1
4.1.3. Potential Harm of ‘What’-type Errors
In the high error condition group, what-type errors are incorrect descriptions of what the AV is doing. To understand the impact of the potential harm of these errors, we conducted a secondary analysis comparing errors that would result in an accident if acted upon by the AV versus those that would not. To test this case, we use LME models with fixed effects for the categorization of potential harm (0 or 1) and random effects for driving scenario and individual differences by participant. This analysis is conducted only on data from the “high” condition in isolation.
We find significant effects for satisfaction with an explanation ($\beta$ = -0.46, t(22) = -2.6, p ¡ .05) and confidence in the AV’s driving ability ($\beta$ = -0.45, t(22) = -2.7, p ¡ .05). Comfort relying on the AV showed results trending towards significance ($\beta$ = -0.38, t(22) = -1.9, p = 0.07). No significant results were found for the reliance preference measure. These results imply that the content of the error – in this case, the potential harm that could result from the error – may impact the effect of the error on outcomes. Specifically, we find evidence that higher gravity errors may have a greater negative impact on some outcomes.
4.2. Impact of Driving Context: Perceived Harm and Difficulty of Driving
4.2.1. Relationship Between Harm and Difficulty
By examining the impact of harm and difficulty on outcome ratings, we can derive an understanding of how these contextual factors influenced our results. Unsurprisingly, we found a strong, positive relationship between the perceived harm and the difficulty of driving in a particular scenario (r = 0.69, p ¡ .001; Adj R2 = .48). Figure 3 shows difficulty and harm ratings by scenario.
Figure 3. Mean Harm and Difficulty Across All Scenarios. Results show harm and difficulty are generally correlated, but differ depending on the specific scenario. We find that these contextual characteristics directly impacted outcome ratings as well as influenced the relationship between errors and outcomes.
[…]
A visualization of the mean Harm and Difficulty scores for all 24 scenarios, depicting the mean Harm and Difficulty ratings for each scenario. Results show harm and difficulty are generally correlated, but differ depending on the specific scenario. We find that these contextual characteristics directly impacted outcome ratings as well as influenced the relationship between errors and outcomes.
4.2.2. Impact of Harm and Difficulty On Main Outcome Ratings
We assess the impact of harm and difficulty on main outcome ratings using LME models with added fixed effects for the average harm and difficulty of scenarios. We maintain the fixed effect of outcome condition as well as random effects for scenario and participant. Using this model, main effects for difficulty and harm give us an understanding of how these factors impacted outcome judgments independent of error condition. Estimates ($\beta$) are given relative to the “accurate” condition. We find that harm significantly impacted comfort relying on the AV ($\beta$ = -0.43, t(38) = -3.7, p ¡ .001), reliance decision ($\beta$ = -0.74, t(62) = -4.4, p ¡ .001), and confidence in the AV’s driving ability ($\beta$ = -0.22, t(41) = -2.2, p ¡ .05). No significant effects were found between harm and explanation satisfaction. We found that difficulty impacted only reliance decision ($\beta$ = 0.60, t(61) = 3.0, p ¡ .01). Interestingly, this was in the opposite direction as harm, where greater difficulty was associated with increased reliance.
The implication is that contextual factors like harm and difficulty affect interaction with an AV system regardless of error condition, with harm potentially being the more influential factor. Specifically, higher harm is generally associated with lower comfort, reliance, and confidence, while higher difficulty is generally associated with higher reliance ratings. There does not appear to be an overall association between harm or difficulty and explanation satisfaction.
4.2.3. Impact Of Harm and Difficulty On Differences Between Error Conditions (Interaction Effects)
Using a similar LME model, we can assess if difficulty or harm impacted the relationship between error level and outcomes by looking at interaction effects. Specifically, we examine if the differences found between each error condition are predicted by a scenario’s average difficulty or harm. In this case, we use the accurate group as a common reference (model intercept) and then compare if the changes of each main outcome from accurate to low and accurate to high vary with respect to difficulty and harm. We find significant interaction effects for several of our outcomes, implying that contextual characteristics like harm and difficulty may be moderating how much of an impact an error may have. Table 5 shows interaction effects for harm, and Table 6 shows interaction effects for difficulty.
Comfort Relying on AV – Starting with comfort relying on the AV, we find that scenario difficulty has more of an impact on comfort in the low error condition than in the accurate condition. Specifically, when difficulty increased, comfort decreased significantly more in the low condition compared to the accurate condition. A similar but opposite effect was found with harm: in the low error condition, comfort increases significantly more with higher harm than in the accurate condition. The implication is that comfort judgments may be more sensitive to the difficulty and harm of the scenario when there are some errors (low condition) compared to when there are no errors (accurate). We do not see difficulty or harm as more impactful on comfort judgments in the high error condition compared to the accurate condition. This implies that these judgments of comfort are likely based on the high magnitude of error (main effect) for the high group, as opposed to being influenced by the difficulty or harm of the driving context (interaction effect) in these cases.
Reliance Decision – The case is similar for reliance decision, though the effects are trending towards significance as opposed to being statistically significant. We find that scenario difficulty has more of an impact on reliance in the low error condition than it did in the accurate condition. Specifically, when difficulty increased, reliance decreased more in the low condition compared to the accurate condition. For harm in the low error condition, higher harm increases reliance significantly more than in the accurate condition. The implication is that reliance judgments may be more sensitive to the difficulty and harm of the scenario when there are some errors (low condition) compared to when there are no errors (accurate). As with comfort, we do not find difficulty or harm as more impactful on reliance in the high error condition compared to the accurate condition. This implies that judgments of reliance when errors are high are likely based on the error itself rather than difficulty or harm.
Satisfaction w/ Expl. – The interaction effects change for satisfaction with an explanation. We do not see difficulty or harm as more impactful on satisfaction in the low error condition compared to the accurate condition, implying that differences in judgments of the explanation are likely based on the explanation quality itself without being influenced by the difficulty or harm of the situation in these cases. The same effect is seen when comparing differences between the accurate and high condition groups for harm. Surprisingly, we find that difficulty differentially impacted explanation satisfaction when in the high condition compared to the accurate condition. Specifically, when difficulty increased, satisfaction increased significantly more when in the high condition than when in the accurate condition. These results together imply that, in most cases, judgments of an explanation are likely based on the explanation quality itself, however, when explanation quality is exceptionally poor and driving is difficult, participants may actually be more forgiving with their judgments.
Confidence in Driving – Lastly, for confidence in the AV’s driving ability, we find a similar trend as comfort and reliance decision. We find that scenario difficulty has more of an impact on confidence in the low error condition than it did in the accurate condition. When difficulty increased, confidence decreased more in the low condition compared to the accurate condition (trending towards significance). For harm in the low error condition, higher harm increases confidence significantly more than in the accurate condition. The implication is that confidence judgments may be more sensitive to the difficulty and harm of the scenario when there are some errors (low condition) compared to when there are no errors (accurate). As with comfort and reliance, we do not find difficulty or harm as more impactful on confidence in the high error condition compared to the accurate condition. This implies that judgments of confidence are more likely based on the main effect of error itself, rather than difficulty or harm.
Table 5. Impact of Harm on Error Condition Differences (LME Models). These show if Harm influences the impact of an error. We find significant effects for several of our outcomes, implying that Harm may moderate how much of an impact an error may have.
Acc.-Low Acc.-High
$\beta$ df t sig. $\beta$ df t sig.
Comfort Relying on AV 0.43 5108 4.2 *** 0.10 5108 0.9 NS
Reliance Decision 0.35 5106 1.8 . 0.19 5106 1.0 NS
Satisfaction w/ Expl. 0.13 5108 1.1 NS -0.15 5108 -1.3 NS
Confidence in Driving 0.27 5108 2.8 ** -0.01 5108 -0.1 NS
Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1
Table 6. Impact of Difficulty on Error Condition Differences (LME Models). These show if Difficulty influences the impact of an error. We find significant effects for several of our outcomes, implying that Difficulty may moderate how much of an impact an error may have.
Acc.-Low Acc.-High
$\beta$ df t sig. $\beta$ df t sig.
Comfort Relying on AV -0.26 5109 -2.1 * 0.1 5109 0.7 NS
Reliance Decision -0.42 5107 -1.9 . -0.16 5107 -0.7 NS
Satisfaction w/ Expl. -0.08 5109 -0.6 NS 0.30 5109 2.1 *
Confidence in Driving -0.21 5108 -1.9 . 0.04 5109 0.3 NS
Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1
4.3. Trust and Expertise
4.3.1. Impact Of Exposure To Errors On Trust
We measured participant trust before and after the experiment was completed to see if exposure to repeated AV explanation mistakes affected AV trust levels as a disposition. In general, using a paired-samples t-test, we did not find that exposure to errors in the experiment impacted trust. Only one result was significant pre-post experiment: participants felt that they “understand why an autonomous vehicle makes decisions” worse after the experiment as compared to before (p ¡ 0.001).
4.3.2. Correlations of Initial Trust and Subjective Expertise with Outcomes
Calculating correlations between initial trust and self-reported expertise level and a participants’ average outcome ratings allows us to assess if trust or expertise can predict how a person will evaluate the AV. We found a strong positive correlation between initial trust and expertise, (r = 0.58, p ¡ 0.001), implying that those who know more about AVs trust them more.
We found moderate positive correlations between expertise and comfort (r = 0.38, p ¡ 0.001), satisfaction (r = 0.35, p ¡ 0.001), and confidence (r = 0.36, p ¡ 0.001), meaning that those with more expertise reported more positively to these outcomes on average. We found high correlations between initial trust and comfort (r = 0.52, p ¡ 0.001) as well as confidence (r = 0.51, p ¡ 0.001). We found moderate correlations between initial trust and reliance (r = 0.30, p ¡ 0.001) as well as satisfaction (r = 0.49, p ¡ 0.001). These effects are consistent with linear mixed-effects models looking at the main effects of expertise and initial trust on outcomes. Together, these results imply that participants with higher initial trust or expertise tended to rate outcomes higher on average.
4.4. Self-Reported Rationale For Trust and Reliance Decisions
4.4.1. Factors Contributing to Rating Decisions
We explicitly inquired upon the basis of reliance decisions by asking participants to rate the relative importance of several factors on their reliance ratings (Figure 4). This was done after the rating task concluded. As a whole, we found that the accuracy of the explanation on ‘what’ the AV is doing was the most important consideration, followed by the explanation accuracy of ‘why’ the AV is doing it, the harm of the situation, the AV’s driving ability, and the difficulty of the driving situation. Prior scenes viewed and prior knowledge did not have much reported impact, though this is unsurprising, as these implicit judgments may be difficult for participants to self-describe. Many of these differences were significant based on ANOVA. Notably, harm was significantly more important than difficulty, while explanation accuracy (both ‘what’ and ‘why’ components) were significantly more important than AV driving ability. This latter effect may be in part due to the salience of these factors based on the study’s manipulation. We did not see significant differences between the importance of ‘what’ and ‘why’ accuracy.
Figure 4. Relative Importance of Factors on Reliance Decisions. This shows the mean importance rating of each factor with respect to a person’s reliance decision. Results show many of the factors of interest in this study were of high importance.
[…]
A visualization of relative importance (mean rating) of seven factors on reliance decisions. Results show many of the factors of interest in this study were of high importance.
4.4.2. Feedback from Participants
Participants were given the opportunity to express their general views on autonomous vehicles on how explanation errors impacted their decision-making specifically. Unsurprisingly, many participants reflected general concerns with AVs, including doubt that they can “adapt to any circumstance … [causing] a bigger accident.” One participant commented that “after a lifetime of driving myself, I’m not sure if I feel comfortable giving over control to a computer.” Desire for control was brought up multiple times. The common sentiment was that giving control to AVs is considered risky.
Regarding exposure to explanation errors, one participant commented that they were “less confident in the reliability of [autonomous] vehicles” after exposure to errors, while another commented that when explanations “were wrong, [it made] my confidence in the system shaky.” These comments reiterate our general study findings on the negative impact of errors on reliance, as well as the holistic judgment of an AV’s ability in general. Some participants broke their decision-making down in terms of individual factor priority. For instance, “if the AV described the action it was taking incorrectly I was a lot less confident and consistently chose to take control myself … if the reason was incorrect, it still impacted my confidence level but not as much.” This reflects the general trend found in our individual factor analysis suggesting ‘what’ information may be more important than ‘why’ rationale, even if this trend was not found to be statically significant.
Regarding the impact of context, a participant noted, “I chose to take control myself in higher risk situations even if the AV did a good job navigating.” This supports this study’s overall findings on the importance of contextual factors like harm on decision-making.
5. Discussion
This study aimed to assess the impact of autonomous vehicle (AV) explanation errors and driving context characteristics on participant comfort relying on an AV, preference to rely on the AV instead of taking control themselves, satisfaction with the explanation, and confidence in the AV’s driving ability. Using a mixed-methods approach with a heterogeneous sample of participants (n = 232), we found that explanation errors and contextual characteristics like driving difficulty and perceived harm have a large impact on how a person may think, feel, and behave towards AVs. Echoing prior work by Nourani et al. (2020), important consideration must also be given to personal factors like initial trust and expertise, which may further impact how a person interacts with the system. Our results provide insight into how to design effective human-AV interactions and interactions with AI-based systems more generally. We discuss key findings and their implications for future AV design in these contexts.
Autonomous vehicle (AV) explanation errors had a detrimental effect on all outcomes. Our results suggest that errors significantly reduced participant comfort relying on an AV, preference to rely on the AV instead of taking control themselves, and satisfaction with the explanation. These effects are unsurprising, as we would expect outcomes related to a person’s reliance decision or satisfaction with an explanation to reflect the system’s performance (Hoff and Bashir, 2015).
We were surprised, however, to find crossover effects between the explanatory performance of the AV and a person’s confidence in the AV’s driving ability, given the AV’s actual demonstrated driving performance remained consistent across all conditions. This crossover effect alludes to the mental model of potential AV users, where evaluation of the explanatory performance of the vehicle and the driving performance of the vehicle are connected. Though connecting explanatory and driving performance may be an intuitive and correct assumption, empirically observing this crossover provides evidence supporting the importance of high quality explanatory interfaces, particularly within safety-critical systems. For the case of AVs, this means that – even if the AV’s driving performance is perfect – if the explanations produced by the AV are not accurate as well, people may still refuse to adopt AV technology. This insight likely generalizes to explanatory communications for other AI-based systems.
The negative impact of errors increased with error magnitude and the potential harm of the error. Expanding on the effect of errors observed in our study, we find that the impact was commensurate with the magnitude of the errors presented. In the study presented here, ‘what’ and ’why’ errors (high condition) in combination had worse outcomes than ‘why’-only errors (low condition). On the surface, this is unsurprising, as more errors demonstrates lower system performance, and these results could simply be the cumulative effects of seeing errors on two parts of the explanation instead of just one.
There remains a high probability that ‘what’ and ‘why’ information errors play distinct roles in the evaluation of the system’s performance, however. This would align with past work showing how description of action and description of justification may differentially impact the way a person interacts with autonomous systems (Koo et al., 2015; Kaufman et al., 2024a; Miller, 2019). When asked explicitly about the factors that contributed to their reliance decisions, participants reported that explanation accuracy (both ‘what’ and ‘why’ components) were both important. Qualitative assessment of participant feedback, however, suggests that ‘what’ information may be more important for reliance outcomes than ‘why’ rationale. This may be due to the potential harm that improper action (‘what’) can have on driving safety, in contrast to improper justification (‘why’). Concretely, a car incorrectly turning right into traffic can cause physical harm, but incorrectly justifying why it is turning right can not. We note that conclusively differentiating the impact of ‘what’ from ‘why’ errors independently will be left for future work.
Supporting the hypothesis that the implications of an error matter in addition to the mere presence of an error, we find that ‘what’ errors – with more dire implications for potential harm – had a greater negative impact on comfort, confidence, and explanation satisfaction. Specifically, in cases when ‘what’ errors would have resulted in vehicle crashes if the AV had acted upon them, people were far less comfortable relying on the AV, had less confidence in the AV’s driving, and preferred the explanation less. We posit that the only reason we did not see differences in reliance preference as well was because the presence of an error altogether already reduced reliance preference to near-minimum levels.
Together, these results suggest that the implications of an explanation error in terms of what they may mean about the system’s performance and consequences of harm, affect how a person thinks, feels, and behaves with an AV. Implications may be implicitly calculated based on the amount and type of error, as well as how these intersect with the external driving context.
Even with accurate explanations, Comfort, Reliance Preference, and Confidence in the AV’s driving ability were fairly low. In the present study, we were primarily interested in the relative differences in outcome ratings based on errors, contextual factors, or individual differences like trust or expertise level. We note, however, that even in cases when the explanation presented was accurate, comfort, reliance, and confidence scores still hovered near the middle of the provided scale. Though the middling scores may be reflective of the overall challenging nature of the driving situations presented in the study, even the most straightforward of the scenarios presented did not have a reliance preference score greater than mid-range. These results reflect the trend found in a plethora of prior work supporting the premise that people do not generally trust or wish to adopt autonomous vehicles.
Contextual factors like the difficulty of driving in a specific situation and the perceived harm of that situation directly affect how people think, feel, and prefer to behave with AVs. We found evidence that driving difficulty and perceived harm directly influenced outcomes regardless of error condition. Specifically, higher contextual harm was associated with lower comfort, reliance, and confidence ratings, while higher difficulty was associated with higher reliance ratings. We did not find direct associations between harm or difficulty and explanation satisfaction.
We attribute the negative impact of harm to the common concern that AVs may malfunction, and a malfunction in a more harmful driving situation may have more severe implications. This reflects the general sentiment that many people think they can drive better than an AV in most situations. The seemingly positive impact of difficulty on reliance may reflect that, in very high difficulty situations, people lack the confidence that they could perform better than the AV and, as such, prefer to rely on the AV in these cases. Indeed, the driving demonstrated by the AV in even the most difficult driving situations presented in the study was high quality.
Explicitly, participants reported contextual harm as significantly more important than driving difficulty for their reliance decision. From a theoretical lens, a person’s concern with difficult driving may stem from the increased likelihood of harm in challenging situations, particularly when they may have concerns with an AV’s ability to perform. This leads us to hypothesize that an overall evaluation of potential for harm is the underlying factor influencing reliance decisions overall. By this lens, difficulty and AV driving performance matter because harm matters. Thus, we hypothesize that harm itself may be the underlying driver of reliance decisions overall.
A person’s decision to rely on an AV (and other outcomes) may be a function of the demonstrated performance of the vehicle – compounded by an implicit evaluation of the overall harm of the situation (in part as a function of difficulty) – weighted against that person’s confidence in their own ability to perform better. This is similar to the model proposed by Hoff and Bashir (2015), with added details on the nested relationship between driving difficulty and harm.
The general implication of these findings is that there is a complex relationship between contextual factors like harm and difficulty, and outcomes like comfort, reliance, and confidence. It is clear that models of AV behavior incorporating internal and external characteristics of the driving situation can bring further insight into how people will interact with AVs (Kaufman et al., 2024b), and future research should examine how these can be best supported by explanatory communication.
Contextual factors moderate the relationship between errors and outcomes in complex, sometimes counter-intuitive ways. By examining the interaction between contextual characteristics and the effects of error conditions, we found that – in some cases – the driving context may influence the amount of impact an error has. The general trend we observed for comfort relying on the AV, reliance preference, and confidence in the AV’s driving ability was that when difficulty increased, outcome scores decreased more in the low error condition compared to the accurate condition, with no differences seen between the accurate and high conditions. The same but opposite effect was found for harm: higher harm increased outcomes effect in the low condition more than the accurate condition, with no difference between the accurate and high condition.
The lack of difference found in the way difficulty and harm moderate the relationship between the accurate and high conditions implies that differences in these judgments are likely based solely on the error level as opposed to difficulty or harm of the driving context. Concretely, the effect of error likely overrides the effect of context when errors are at the extremes. When errors are in the middle – such as for the low error condition – the impact of the error appears to be more strongly moderated by the context. This may be because the ramifications of ‘why’-type errors are not strong enough to drive effects fully based on the presence of the error, and instead the amount of change that the error can cause is based on how dire the implications may be in a particular context. The direction of effect, in this case, is similar to the main effects explained previously.
For explanation satisfaction, we do not see difficulty or harm as more impactful on satisfaction in the low error condition, if we compare this to the accurate condition. This implies that differences in satisfaction are likely based on the explanation quality itself in these cases. The same effect is seen between the accurate and high condition for harm, but not for difficulty. Surprisingly, we found that when difficulty increased, satisfaction increased significantly more when in the high condition than when in the accurate condition. Together, these results imply that, in most cases, judgments of an explanation are based primarily on the explanation quality itself, however, when explanation quality is especially poor and driving is difficult, participants may actually be more forgiving of the AV.
Taken together, results from this analysis on interaction effects illustrate a nuanced and often complex relationship between driving difficulty, perceived harm and error condition when predicting our main outcomes. This nuance alludes to a prioritization of how different aspects of the driving situation – including harm, difficulty, and explanation errors – combine to form judgments of comfort in the AV, reliance preference, satisfaction with an explanation, and confidence in an AV’s driving ability.
Trust and expertise influenced how people think, feel, and behave towards the AV. As discussed, how a person thinks and behaves towards AVs is not just a function of their external environment. We find evidence that internal characteristics, such as a participant’s trust and expertise, may also have played a role in the study’s results. Participants with higher initial trust or expertise tended to rate study outcomes higher on average. We also found that those with higher AV expertise trusted AVs more in general. As with the function on driving reliance described previously, these findings also align with Hoff and Bashir’s model of trust with AI systems, where a person’s prior experience impacts their reliance with the system (Hoff and Bashir, 2015). Those with higher expertise may also have other underlying characteristics, like an affinity for new technologies, which may have influenced their trust. We attribute the finding that priors were not reported as explicitly important for reliance decisions in our post-evaluation to the common difficulty articulating the impact of implicit dispositions.
Though trust may be influential, we did not find evidence that exposure to AV mistakes due to the study’s manipulation had a large effect on a person’s overall trust level. A single exception was that participants reported less understanding of why AVs make decisions once the study concluded. This makes sense given the inconsistent explanations for behavior they had received, which may have impacted a person’s mental model of AV decision-making.
5.1. Implications For Autonomous Vehicle Design and Research
Our results have important implications for future AV design and research. Understanding the consequences of AV explanation errors and contextual characteristics like driving difficulty and perceived harm are necessary first steps towards designing vehicles which may be more trustworthy, reliable, and satisfactory for people interacting with the AV. In this section, we will discuss how our study’s insights can be used to design more trustworthy AVs.
The foremost implication of this work is to emphasize the importance of designing systems that can produce accurate explanations for AV decisions. This implication is not too surprising, as why would designers ever intentionally produce inaccurate explanations? It becomes more meaningful, however, when testing explanatory systems before deployment. Our results indicate that even with a well-functioning driving system, if explanations are not of high quality, people still won’t want to use the AV.
The crossover between a person’s evaluation of explanation performance and driving performance may be of particular interest to AV designers. If this crossover is a concern – as it should be in the case of deployment – the most obvious solution would be to work on improving explanations so that they are on par with driving ability. In the meantime, specific UI features may be implemented to help users separate their evaluations of explanations from driving performance, such as by presenting distinct indicators for each system reassuring a rider that even if the explanation is messed up, the car can still drive sufficiently. This may be challenging or misleading, however, particularly if the explanatory system and driving system are indeed connected.
Though not tested in the present study, our results also have implications for conditionally automated vehicles – those which may require human takeover in difficult or computationally complex driving situations (SAE Level 3) (Inagaki and Sheridan, 2019; Ayoub et al., 2021a). There is the possibility that explanation errors produced in conditional contexts may result in inappropriate behaviors by human passengers. If users lose trust in the AV due to explanation errors, for example, they may react unpredictably or take over control at inappropriate times, potentially leading to accidents.
For both the fully autonomous and conditionally autonomous cases, our results indicate that explanatory systems should be deployed with confidence in their accuracy or potentially not at all. Future work can estimate the exact outcome differences between no explanation and inaccurate explanations, but combining the results of our study and prior findings on the importance of AI explainability by Gunning and Aha (2019) and Miller (2019) suggests that accurate explanations should be opted for whenever possible, and deploying untested or inconsistent systems can be disastrous.
In terms of design priority, our work suggests that if explanation designers are going to focus on improving a single aspect of the explanation, they should optimize for accurate ‘what’ information before ‘why’ justification. Though both may be necessary in the long run, providing a correct description of behavior may produce a higher value return on effort for teams looking to find a place to start. Just as they have for differentiating the effects of accurate ‘what’ and ‘why’ explanations of driving behavior, future work should delineate the potentially different impacts ‘why’ and ‘what’ errors may have from each other in isolation. This can allow more detailed and conclusive conclusions on their relative importance to outcomes such as those in the present study.
It is important to note that, while explanation satisfaction scores were generally higher than other outcome scores in the accurate condition, explanations were still nowhere near ceiling level. This means that there are likely ways in which our explanations could have improved. Prior work has emphasized outcome differences based on informational content, modality (such as including visualizations or haptic feedback), timing, or interactivity (Miller, 2019; Kaufman et al., 2024a; Avetisyan et al., 2022; Goldman and Bustin, 2022). In this regard, future AV research and design teams should continue iterating on these aspects of explanation design and human-machine interface (HMI) development in order to fine-tune explainable systems.
Another major implication of this work is to orient AV designers to the potentially different outcomes that may result from interactions by people with different traits or while operating in driving environments. Specifically, context-aware or personalized explanations may be necessary in cases where the driving difficulty or perceived harm are classified as greater, or for people with little expertise or initial trust. It is likely that designers will need to focus on limitations on cognitive resources like attention and load, particularly for high difficulty scenarios. Other personal traits like risk preferences or personality may also be a basis for personalization (Böckle et al., 2021). Segmenting populations or isolating particular driving conditions to test future explanation designs would be an effective method to figure out how to meet the differing needs of diverse user groups (Kaufman et al., 2024b).
Taken in full, our results provide insight into the impact of explanation errors, and provide direction for future AV researchers and designers to improve human-AV communication. Given the serious negative repercussions that may result from errors, our results also provide a foundation for the creation of ethical or regulatory guidelines which dictate the testing and accuracy of AV explanations before they can be deployed to real-world consumers.
5.2. Limitations and Future Study
As with any study, this research is not without limitations. First, though study findings were generally robust and fit within logical narratives, there is the possibility that findings resulting from online data collection and based on simulations may not generalize to real-world attitudes or behavior. A large effort was put into making driving simulations as realistic as possible, however, we could not test the effect of explanation errors on real roads out of concern for participant safety. Future work should seek to test the impact of errors in a real-world environment. A similar concern may be found for study outcomes: though results on reported reliance or comfort may be clear in a research context, there is no guarantee that explicit declarations of thought or behavior will remain consistent when immersed in a real-world driving context with real consequences of bodily or financial harm. This is a concern for all studies which rely on survey measures or explicit participant statements as a proxy for real-world behavior. It is possible that seeing each scenario three times (with different explanations) may have impacted the ratings provided. This could be mitigated in the future by showing participants only one video per scenario, randomized by error condition. Finally, we examined proximal explanations for action (“braking”) and cause of action (“… a pedestrian is crossing the road.”) presented in written fashion and auditory fashion. Future work can examine explanations of distal causes (“pedestrian in the road … because there is an obstacle on the sidewalk”) which could provide additional context for why a driving situation is happening in the first place. These could be explained using potentially different modalities of presentation.
6. Conclusion
In a simulated driving study with 232 participants we tested how autonomous vehicle (AV) explanation errors, driving context characteristics (perceived harm and driving difficulty), and personal traits (prior trust and expertise) impact four driving perception and behavior-related outcomes: comfort relying on an AV, preference to rely on the AV instead of taking control themselves, satisfaction with the explanation, and confidence in the AV’s driving ability.
Our results indicate that explanation errors, contextual characteristics, and personal traits have a large impact on how a person may think, feel, and behave towards AVs. Explanation errors negatively affected all outcomes. Surprisingly, this included reduced ratings of the AV’s driving ability, despite driving performance remaining constant. The negative impact of errors increased with error magnitude and the potential harm of the error, providing evidence that implications of an error matter in addition to the mere presence of an error. Harm and driving difficulty directly impacted outcomes as well as moderated the relationship between errors and outcomes. In general, harm was associated with lower comfort, reliance, and confidence ratings, while driving difficulty was associated with higher reliance ratings. Overall harm was the more important contextual factor of consideration. In terms of a decision function for AV reliance, perceived harm – influenced by driving difficulty – may underlie the evaluation between a person’s confidence in the AV’s performance compared to their own ability. We found that individuals with higher expertise tended to trust AVs more, and these each correlated with more positive outcome ratings in turn.
Overall, our results emphasize the need for accurate and contextually adaptive AV explanations to foster trust, reliance, satisfaction, and confidence. Understanding the ramifications of explanation errors can help future AV research and design teams prioritize design and better understand the impacts of their design choices. They also provide a foundation for context-aware design, personalized explanation interfaces, and potential ethical or regulatory guidelines for the deployment of explainable AI (XAI) systems for autonomous vehicles that are safe and trustworthy.
Acknowledgements.
We would like to thank S. Sapre, R. Bhide, C. Lee, J. Molina, and E. Lee for simulator development and testing.